OpenStudy (anonymous):

answer for cookies! f(t) = tan(e^3t) + e^tan 3t

zepdrix (zepdrix):

Oooo cookies :OOO

zepdrix (zepdrix):

What are we going? Differentiating?

OpenStudy (anonymous):

yes sir

zepdrix (zepdrix):

Lady bunch! I thought you already did this problem a few days ago! >.< Maybe I'm mistaken though...

OpenStudy (anonymous):

ran out of cookies, had to do it again

zepdrix (zepdrix):

\[\Large\rm f(t) = \tan(e^{3t}) + e^{\tan 3t}\]

zepdrix (zepdrix):

Do you remember your derivatives for e^x and tan x?

OpenStudy (anonymous):

e^x equal itself tanx is sec^2x

zepdrix (zepdrix):

Mmm k good good.

zepdrix (zepdrix):

Sorry power went out :c Did I lose my cookie. Ok ok we're still good.. sec

zepdrix (zepdrix):

So using your knowledge of those derivatives, it tells us that when we differentiate we should get something like: \[\Large\rm f'(t) = \sec^2(e^{3t}) + e^{\tan 3t}\] But we need to apply our chain rules, yes?

zepdrix (zepdrix):

\[\Large\rm f'(t) = \sec^2(e^{3t})\color{royalblue}{\left(e^{3t}\right)'}+ e^{\tan 3t}\color{royalblue}{\left(\tan 3t\right)'}\]So we need to multiply by the derivative of the inner functions.

zepdrix (zepdrix):

So again, as you said before, our exponential gives us the same thing back:\[\Large\rm f'(t) = \sec^2(e^{3t})\color{orangered}{\left(e^{3t}\right)}\color{royalblue}{(3t)'}+ e^{\tan 3t}\color{royalblue}{\left(\tan 3t\right)'}\]But we have to chain rule again!! :O

zepdrix (zepdrix):

\[\Large\rm f'(t) = \sec^2(e^{3t})\color{orangered}{\left(e^{3t}\right)}\color{orangered}{(3)}+ e^{\tan 3t}\color{royalblue}{\left(\tan 3t\right)'}\]

zepdrix (zepdrix):

So we've fully differentiated the first term. How bout that other term, Baby Bump? What do you think? :) Give it a try! What should we get?

zepdrix (zepdrix):

Comeon Banana Bunch! Don't give up on me now >:U Participateeeee

OpenStudy (anonymous):

\[\sec^23t(3)\]

zepdrix (zepdrix):

Yay good job \c:/